OpenStudy (anonymous):

(x^4)^4*(x^3)^5 simplified would be x^33 is this correct?

OpenStudy (dumbcow):

no it would be 31 4*4=16 3*5=15 16+15=31

OpenStudy (anonymous):

ok but x*x is x^2 so wouldn't you add two to that?

OpenStudy (theviper):

But here it is x^16 * x^15

OpenStudy (theviper):

Got it:)

OpenStudy (anonymous):

ok yea I think I get it thanks idk y i wanted to add the 2 in there

OpenStudy (anonymous):

Remember the rules. \[x^{2} * x^{3} = x^{2 + 3} = x^{5}\] \[(x^{2})^{3} = x^{2 * 3} = x^{6}\]

OpenStudy (theviper):

x*x can be x^2 but here it is x^16 * x^15

OpenStudy (theviper):

Any confusion?????

OpenStudy (anonymous):

nope pretty sure it was just a blonde moment...thank you

OpenStudy (theviper):

U r welcome!!!!!